OpenStudy (anonymous):

if f(x)=x^2-6 and g(x)=2^x-1 determine the value of (gof)(-3)

OpenStudy (mertsj):

You could find gof and plug in-3 or you could find f(-3) and plug that answer into g. Which do you want to do?

OpenStudy (anonymous):

how would you find the gof?

OpenStudy (anonymous):

would you plug the formula for g into f? or vice versa?

OpenStudy (anonymous):

g(f(-3)) = g((-3)^2 - 6) = g(9-6) = g(3) = 2*3 - 1 = 6 - 1 = 5

OpenStudy (anonymous):

f(-3)=(-3)^2-6=3 so g(3)=2(3)-1=5

OpenStudy (anonymous):

(gof)(-3) = 5

OpenStudy (anonymous):

thank you so much!